### v5

# Statistical Inference of Constrained Stochastic Optimization via Sketched Sequential Quadratic Programming

Abstract
We consider online statistical inference of constrained stochastic nonlinear optimization problems. We apply the Stochastic Sequential Quadratic Programming (StoSQP) method to solve these problems, which can be regarded as applying second-order Newton’s method to the Karush-Kuhn-Tucker (KKT) conditions. In each iteration, the StoSQP method computes the Newton direction by solving a quadratic program, and then selects a proper adaptive stepsize ${\bar{\alpha}}_{t}$ to update the primal-dual iterate. To reduce dominant computational cost of the method, we inexactly solve the quadratic program in each iteration by employing an iterative sketching solver. Notably, the approximation error of the sketching solver need not vanish as iterations proceed, meaning that the per-iteration computational cost does not blow up. For the above StoSQP method, we show that under mild assumptions, the rescaled primal-dual sequence $1/\sqrt{{\bar{\alpha}}_{t}}\cdot({\bm{x}}_{t}-{\bm{x}}^{\star},{\bm{\lambda}}_{t}-{\bm{\lambda}}^{\star})$ converges to a mean-zero Gaussian distribution with a nontrivial covariance matrix depending on the underlying sketching distribution. To perform inference in practice, we also analyze a plug-in covariance matrix estimator. We illustrate the asymptotic normality result of the method both on benchmark nonlinear problems in CUTEst test set and on linearly/nonlinearly constrained regression problems.
Keywords: constrained stochastic optimization, Newton sketching, online inference, uncertainty quantification, randomized numerical linear algebra
1 Introduction
We consider equality-constrained stochastic nonlinear optimization problems of the form
$\min_{{\bm{x}}\in\mathbb{R}^{d}}\;\;f({\bm{x}})=\mathbb{E}_{{\mathcal{P}}}[F({\bm{x}};\xi)],\quad\text{s.t.}\;\;c({\bm{x}})={\bm{0}},$ (1.1)
[…]
In this paper, we are particularly interested in performing statistical inference on a (local) primal-dual solution $({\bm{x}}^{\star},{\bm{\lambda}}^{\star})$ of Problem (1.1). To that end, the classical (offline) approach often generates $N$ samples $\xi_{1},\ldots,\xi_{N}\sim{\mathcal{P}}$ iid, and then solves the corresponding empirical risk minimization (ERM) problem:
$\min_{{\bm{x}}\in\mathbb{R}^{d}}\;\;\widehat{f}({\bm{x}})=\frac{1}{N}\sum_{i=1}^{N}F({\bm{x}};\xi_{i}),\quad\text{s.t.}\;\;c({\bm{x}})={\bm{0}}.$ (1.2)
Under certain regularity conditions, we can establish the asymptotic consistency and normality of the minimizer $(\widehat{{\bm{x}}}_{N},\widehat{{\bm{\lambda}}}_{N})$ of (1.2), also called constrained $M$-estimator, given by
$\footnotesize\sqrt{N}\begin{pmatrix}\widehat{{\bm{x}}}_{N}-{\bm{x}}^{\star}\\ \widehat{{\bm{\lambda}}}_{N}-{\bm{\lambda}}^{\star}\end{pmatrix}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\left(\begin{pmatrix}{\bm{0}}\\ {\bm{0}}\end{pmatrix},\begin{pmatrix}\nabla_{{\bm{x}}}^{2}\mathcal{L}^{\star}&(G^{\star})^{T}\\ G^{\star}&{\bm{0}}\end{pmatrix}^{-1}\begin{pmatrix}\text{cov}(\nabla F({\bm{x}}^{\star};\xi))&{\bm{0}}\\ {\bm{0}}&{\bm{0}}\end{pmatrix}\begin{pmatrix}\nabla_{{\bm{x}}}^{2}\mathcal{L}^{\star}&(G^{\star})^{T}\\ G^{\star}&{\bm{0}}\end{pmatrix}^{-1}\right),$ (1.3)
where $\mathcal{L}({\bm{x}},{\bm{\lambda}})=f({\bm{x}})+{\bm{\lambda}}^{T}c({\bm{x}})$ is the Lagrangian function with ${\bm{\lambda}}\in\mathbb{R}^{m}$ being the dual variables associated with the constraints, $\nabla_{{\bm{x}}}^{2}\mathcal{L}^{\star}$ is the Lagrangian Hessian with respect to ${\bm{x}}$ evaluated at $({\bm{x}}^{\star},{\bm{\lambda}}^{\star})$, and $G^{\star}=\nabla c({\bm{x}}^{\star})\in\mathbb{R}^{m\times d}$ is the constraints Jacobian. See (Shapiro2014Lectures, Chapter 5) for the result of (1.3), and Duchi2021Asymptotic and Davis2024Asymptotic for showing (1.3) attains the minimax optimality. Numerous methods can be applied to solve constrained ERM (1.2), including (exact) penalty methods, augmented Lagrangian methods, and sequential quadratic programming (SQP) methods (Nocedal2006Numerical).
Given the prevalence of streaming datasets in modern problems, offline methods that require dealing with a large batch set in each step are less attractive. It is desirable to design fully online methods, where only a single sample is used in each step, and to perform online statistical inference by leveraging those methods. Without constraints, one can apply stochastic gradient descent (SGD) and its many variates, whose statistical properties (e.g., asymptotic normality) have been comprehensively studied from different aspects (Robbins1951stochastic; Kiefer1952Stochastic; Polyak1992Acceleration; Ruppert1988Efficient). However, unlike solving unconstrained stochastic programs, there are limited methods proposed for constrained stochastic programs (1.1) that enable online statistical inference. We refer to Section 2.2 for a detailed literature review. One potential exception is the projection-based SGD recently studied in Duchi2021Asymptotic; Davis2024Asymptotic. Although the literature has shown that projected methods also exhibit asymptotic normality, there are two major concerns when applying these methods for practical statistical inference.
  1. (a)
    It is unclear how to online estimate the limiting covariance matrix based on the projected primal iterates. Due to the absence of dual update, the joint primal-dual normality as in (1.3) is not (at least, immediately) achievable for projected methods. Even for the primal normality, the covariance matrix still depends on the dual solution through the Lagrangian Hessian (cf. (1.3)). However, due to intrinsic objective noise, simply using primal iterates and optimality conditions to solve for the dual solution does not yield a consistent dual estimator for the underlying plug-in covariance estimation. One possible resolution is to draw inspiration from long-run variance estimations of stationary processes, and design batch-means covariance estimators that utilize only the projected primal iterates themselves. That said, this approach is highly nontrivial for projected methods (because of the non-stationarity), as studied in the context of vanilla SGD methods (Chen2020Statistical; Zhu2021Online).
  2. (b)
    There are prevalent scenarios where the projection operator becomes intractable. For example, when the constraint function $c({\bm{x}})$ is nonlinear and nonconvex, as in physics-informed neural networks (cf. Section 2.1), the projection onto the manifold $\{{\bm{x}}\in\mathbb{R}^{d}:c({\bm{x}})={\bm{0}}\}$ is generally intractable. Additionally, if we only have local information about the constraint function $c({\bm{x}})$ (e.g., function evaluation and Jacobian) at any given point ${\bm{x}}$, as in CUTEst benchmark nonlinear problems (cf. Section LABEL:sec:5), the projection operator is not computable either, which requires a global characterization of the constraint set.
To perform online statistical inference of Problem (1.1) without relying on projections, we draw inspiration from a recent growing series of literature in numerical optimization, which develops various stochastic sequential quadratic programming (StoSQP) methods for (1.1). The SQP methods can be regarded as second-order Newton’s methods applied to the Karush-Kuhn-Tucker (KKT) conditions. In particular, the StoSQP methods compute a stochastic Newton direction in each iteration by solving a quadratic program, whose objective model is estimated using the new sample. Then, the methods select a proper stepsize to achieve a sufficient reduction on the merit function, which balances the optimality and feasibility of the iterates. We refer to Na2022adaptive; Na2023Inequality; Berahas2021Sequential; Berahas2023Stochastic; Fang2024Fully and references therein for recent StoSQP designs and their promising performance on various problems. The aforementioned literature established the global convergence of StoSQP methods, where the KKT residual $\|\nabla\mathcal{L}({\bm{x}}_{t},{\bm{\lambda}}_{t})\|$ converges to zero almost surely or in expectation for any initialization. However, in contrast to Duchi2021Asymptotic; Davis2024Asymptotic studying projection-based methods, these literature overlooked the statistical properties and failed to quantify the uncertainty inherent in the StoSQP methods, which is yet crucial for applying these methods on online statistical inference tasks. Thus, we pose the following question:
Can we perform online inference on $({\bm{x}}^{\star},{\bm{\lambda}}^{\star})$ based on the StoSQP iterates, while further reducing the computational cost of existing second-order StoSQP methods?
In this paper, we answer this question by complementing the global convergence guarantees and establishing the local asymptotic properties of existing StoSQP methods. Specifically, we focus on an Adaptive Inexact StoSQP scheme, referred to as AI-StoSQP. By adaptive we mean that the scheme inherits the critical merit of numerical StoSQP designs (Berahas2021Sequential; Curtis2021Inexact; Berahas2023Stochastic), allowing for an adaptive stepsize ${\bar{\alpha}}_{t}$ for the Newton direction. In other words, we do not compromise the adaptivity of StoSQP to establish the local convergence guarantees. By inexact we mean that the scheme further reduces the computational cost of StoSQP methods by applying an iterative sketching solver to inexactly solve the Newton system in each step (Strohmer2008Randomized; Gower2015Randomized; Pilanci2016Iterative; Pilanci2017Newton; Lacotte2020Optimal). Solving Newton systems is considered the most computationally expensive step of second-order methods; and randomized solvers offer advantages over deterministic solvers by requiring less flops and memory when equipped with proper sketching matrices (e.g., sparse sketches). Notably, we perform a constant number of sketching steps; thus, the per-iteration computational cost remains fixed even near stationarity.
For the above sketched StoSQP scheme, we quantify its uncertainty consisting of three components: random sampling, random sketching, and random stepsize. We establish the asymptotic normality of the primal-dual iterate:
$1/\sqrt{{\bar{\alpha}}_{t}}\cdot({\bm{x}}_{t}-{\bm{x}}^{\star},{\bm{\lambda}}_{t}-{\bm{\lambda}}^{\star})\stackrel{{\scriptstyle d}}{{\longrightarrow}}{\mathcal{N}}(0,\Xi^{\star}),$ (1.4)
where the limiting covariance $\Xi^{\star}$ solves a Lyapunov equation that depends on the underlying sketching distribution used in the sketching solver. Let ${\Omega^{\star}}$ denote the limiting covariance of constrained $M$-estimator in (1.3). Our result suggests that if ${\bar{\alpha}}_{t}\asymp 1/t$, then we have two cases:
[…]
This implies that (i) if we solve Newton systems exactly, then online StoSQP estimator (even with adaptive stepsizes) achieves the same estimation efficiency as offline $M$-estimator. In fact, if we focus solely on the primal variables ${\bm{x}}$, the marginal covariance of ${\Xi^{\star}}\hskip-1.0pt=\hskip-1.0pt{\Omega^{\star}}$ also matches the limiting covariance of online projection-based estimators established in Duchi2021Asymptotic; Davis2024Asymptotic, which is known to be asymptotic minimax optimal (see Remark LABEL:rem:2). (ii) If we solve Newton systems inexactly, then the sketching solver hurts the asymptotic optimality of StoSQP as ${\Xi^{\star}}\succeq{\Omega^{\star}}$. Fortunately, the hurt is tolerable as seen from the bound (cf. Corollary LABEL:cor:3)
$\|{\Xi^{\star}}-{\Omega^{\star}}\|\lesssim\rho^{\tau}\hskip 8.5359pt\text{for some }\rho\in(0,1),$
where $\tau$ is the number of iterations we run for the sketching solver at each step. In addition to asymptotic normality, we also present some by-product results of independent interest, including the local convergence rate, sample complexity, and the Berry-Esseen bound that quantitatively measures the convergence rate in (1.4). To facilitate practical inference, we also analyze a plug-in covariance estimator that can be computed in online fashion. We illustrate our results on benchmark nonlinear problems in CUTEst test set and on linearly/nonlinearly constrained regression problems.
Structure of the paper. We introduce some motivating examples of Problem (1.1) and provide a literature review in Section 2. Then, we introduce AI-StoSQP in Section 3 and prove the global almost sure convergence with iteration complexity in Section 4. Asymptotic normality with covariance estimation is established in Section LABEL:sec:4. Experiments and conclusions are presented in Sections LABEL:sec:5 and LABEL:sec:6, respectively. We defer all the proofs to the appendices.
[…]
2 Applications and Literature Review
We present two motivating examples of (1.1) in Section 2.1, and then review related literature in Section 2.2.
2.1 Motivating examples
Many statistical and machine learning problems can be cast into the form of Problem (1.1).
Example 2.1 (Constrained regression problems).
Let $\xi_{t}=(\xi_{{\bm{a}}_{t}},\xi_{b_{t}})$ be the $t$-th sample, where $\xi_{{\bm{a}}_{t}}\in\mathbb{R}^{d}$ is the feature vector independently drawn from some multivariate distribution and $\xi_{b_{t}}$ is the response. We consider different regression models, such as
[…]
where ${\bm{x}}^{\star}\in\mathbb{R}^{d}$ is the true model parameter. For the above models, we define the corresponding loss functions at ${\bm{x}}$:
linear models: $F({\bm{x}};\xi_{t})=\frac{1}{2}(\xi_{{\bm{a}}_{t}}^{T}{\bm{x}}-\xi_{b_{t}})^{2},$
logistic models: $F({\bm{x}};\xi_{t})=\log\left(1+\exp(-\xi_{b_{t}}\cdot\xi_{{\bm{a}}_{t}}^{T}{\bm{x}})\right).$
Then, we can verify that ${\bm{x}}^{\star}=\mathop{\mathrm{argmin}}_{{\bm{x}}}f({\bm{x}})=\mathop{\mathrm{argmin}}_{{\bm{x}}}\mathbb{E}[F({\bm{x}};\xi)]$. In many cases, we access prior information about the model parameters, which is encoded as constraints. For example, in portfolio selection, ${\bm{x}}$ represents the portfolio allocation vector that satisfies ${\bm{x}}^{T}{\bm{1}}=1$ (total allocation is 100%). We may also fix the target percentage on each sector or each region, which is translated into the constraint $A{\bm{x}}={\bm{d}}$. See (Fan2007Variable, (4.3, 4.4)), (Fan2012Vast, (2.1)), (Du2022High, (1)) and references therein for such applications. In principle component analysis and semiparametric single/multiple index regressions, we enforce ${\bm{x}}$ to have a unit norm to address the identifiability issue, leading to a nonlinear constraint $\|{\bm{x}}\|^{2}=1$. We point to Kaufman1978Method; Kirkegaard1972POSITRONFIT; Sen1979Asymptotic; Nagaraj1991Estimation; Dupacova1988Asymptotic; Shapiro2000asymptotics; Na2019High; Na2021High for various examples of linearly/nonlinearly constrained estimation problems. For some constrained estimation problems, projecting into the feasible set can be intractable. For instance, in factor analysis, researchers may estimate a covariance matrix $\Sigma$ under so-called tetrad constraints: $\Sigma_{i_{1}i_{2}}\Sigma_{i_{3}i_{4}}-\Sigma_{i_{1}i_{4}}\Sigma_{i_{2}i_{3}}=0$ for every set of four distinct variables $\{i_{1},i_{2},i_{3},i_{4}\}$ (Bollen2000tetrad; Drton2016Wald; Sturma2024Testing). For such highly nonlinear constraints, the linear-quadratic approximation performed in SQP can be a promising resolution.
Example 2.2 (Physics-informed machine learning).
Recent decades have seen machine learning (ML) making significant inroads into science. The major task in ML is to learn an unknown mapping ${\bm{z}}(\cdot):\mathcal{A}\rightarrow\mathcal{B}$ from data that can perform well in the downstream tasks. Since ${\bm{z}}(\cdot)$ is infinite-dimensional, one key step in ML is to use neural networks (NNs) to parameterize ${\bm{z}}(\cdot)$ as ${\bm{z}}_{{\bm{x}}}(\cdot)$, and learn the optimal weight parameters ${\bm{x}}\in\mathbb{R}^{d}$ instead (called function approximation). One of the trending topics in ML now is physics-informed ML, where one requires ${\bm{z}}(\cdot)$ to obey some physical principles that are often characterized by partial differential equations (PDEs) (Karniadakis2021Physics; Cuomo2022Scientific). In such applications, we can use the squared loss function, defined for the $t$-th sample $\xi_{t}=(\xi_{{\bm{a}}_{t}},\xi_{{\bm{b}}_{t}})\in\mathcal{A}\times\mathcal{B}$ as
$F({\bm{x}};\xi_{t})=\frac{1}{2}\left({\bm{z}}_{{\bm{x}}}(\xi_{{\bm{a}}_{t}})-\xi_{{\bm{b}}_{t}}\right)^{2}.$
Here, $\xi_{{\bm{a}}_{t}}$ is NN inputs that can be spatial and/or temporal coordinates; $\xi_{{\bm{b}}_{t}}$ is measurements that can be speed, velocity, and temperature, etc; and ${\bm{z}}_{{\bm{x}}}\in\mathcal{C}^{\infty}(\mathcal{A},\mathcal{B})$ is NN architecture. Let $\mathcal{F}:\mathcal{C}^{\infty}(\mathcal{A},\mathcal{B})\rightarrow\mathcal{C}^{\infty}(\mathcal{A},\mathcal{B})$ be the PDE operator, which encodes the underlying physical law (e.g., energy conservation law). We aim to find optimal weights ${\bm{x}}^{\star}$ that not only minimize the mean squared error of observed data, but also satisfy the constraints $\mathcal{F}({\bm{z}}_{{\bm{x}}})={\bm{0}}$. To this end, we select some leverage points $\{\xi_{{\bm{a}}_{i}}^{\prime}\}_{i=1}^{m}$ in $\mathcal{A}$ and impose deterministic constraints:
$\mathcal{F}({\bm{z}}_{{\bm{x}}})(\xi_{{\bm{a}}_{i}}^{\prime})={\bm{0}},\quad\forall i=1,2,\ldots,m.$
Here, we abuse the notation ${\bm{0}}$ to denote either a zero mapping of $\mathcal{C}^{\infty}(\mathcal{A},\mathcal{B})$ or a zero element of $\mathcal{B}$. For more details on this problem formulation, see (Lu2021Physics, (2.3))(Krishnapriyan2021Characterizing, (2)), and references therein. Due to the nonlinearity nature of NNs, projection-free methods are desired, and SQP can achieve competitive performance compared to penalty methods and augmented Lagrangian methods (Cheng2024Physics).
2.2 Related literature and contribution
There are numerous methods for solving constrained optimization problems, such as projection-based methods, penalty methods, augmented Lagrangian methods, and sequential quadratic programming (SQP) methods (Nocedal2006Numerical). This paper particularly considers solving constrained stochastic optimization problems via Stochastic SQP (StoSQP) methods, which can be regarded as an application of stochastic Newton’s method on constrained problems. Berahas2021Sequential designed the very first online StoSQP scheme. At each step, the method selects a suitable penalty parameter of an $\ell_{1}$-penalized objective; ensures the Newton direction produces a sufficient reduction on the penalized objective; and then selects an adaptive stepsize $\beta_{t}\leq{\bar{\alpha}}_{t}\leq\eta_{t}=\beta_{t}+\chi_{t}$ based on input sequences $\beta_{t}$ and $\chi_{t}=O(\beta_{t}^{2})$. An alternative StoSQP scheme was then reported in Na2022adaptive, where ${\bar{\alpha}}_{t}$ is selected by performing stochastic line search on the augmented Lagrangian with batch sizes increasing as iteration proceeds. Subsequently, Curtis2021Inexact; Na2023Inequality; Berahas2023Stochastic; Fang2024Fully proposed different variates of StoSQP to cope with inequality constraints, degenerate constraints, etc. These works all proved the global convergence of StoSQP methods — the KKT residual $\|\nabla\mathcal{L}({\bm{x}}_{t},{\bm{\lambda}}_{t})\|$ converges to zero from any initialization. However, they fall short of uncertainty quantification and online statistical inference goals.
On the other hand, a growing body of literature leverages optimization procedures to facilitate online inference, starting with Robbins1951stochastic; Kiefer1952Stochastic and continuing through Robbins1971convergence; Fabian1973Asymptotically; Ermoliev1983Stochastic. To study the asymptotic distribution of stochastic gradient descent (SGD), Ruppert1988Efficient and Polyak1992Acceleration averaged SGD iterates and established the optimal central limit theorem rate. Toulis2017Asymptotic designed an implicit SGD method and showed the asymptotics of averaged implicit SGD iterates. Li2018Statistical designed an inference procedure for constant-stepsize SGD by averaging the iterates with recurrent burn-in periods. Mou2020linear further showed the asymptotic covariance of constant-stepsize SGD with Ployak-Ruppert averaging. Liang2019Statistical designed a moment-adjusted SGD method and provided non-asymptotic results that characterize the statistical distribution as the batch size of each step tends to infinity. Chen2020Statistical and Zhu2021Online proposed different covariance matrix estimators constructed by grouping SGD iterates. Additionally, Chen2021First designed a distributed method for the inference of non-differentiable convex problems; Roy2023Online analyzed a batch-mean covariance estimator under a Markovian sampling setup; and Duchi2021Asymptotic and Davis2024Asymptotic applied projection-based SGD methods for the inference of inequality-constrained convex problems. The aforementioned literature all studied first-order methods with deterministic stepsizes.
The asymptotics of second-order Newton’s methods for unconstrained problems have recently been investigated. Bercu2020Efficient designed an online Newton’s method for logistic regression, and Boyer2023asymptotic generalized that method to general regression problems. Compared to first-order methods that often consider averaged iterates and/or exclude the stepsize $1/t$ due to technical challenges, both works showed the normality of the last iterate with $1/t$ stepsize. However, those analyses are not applicable to our study for two reasons. First, they studied unconstrained regression problems with objectives in the form $F({\bm{x}}^{T}\xi)$, resulting in objective Hessians owning rank-one updates that cannot be employed for our general problem (1.1). Second, they solved Newton systems exactly and utilized $1/t$ deterministic stepsize. In contrast, we use a randomized sketching solver to solve Newton systems inexactly to reduce the computational cost associated with higher-order methods, along with an adaptive random stepsize inspired by numerical designs in Berahas2021Sequential. Both of these components affect the uncertainty quantification and lead to a different normality result (cf. (1.5)). To our knowledge, this is the first work that performs online inference by taking into account not only the randomness of samples but also the randomness of computation (i.e., sketching and stepsize); the latter is particularly important for making second-order methods computationally promising.
We briefly review the literature on sketched Newton methods. Compared to the works below, the sketching step in StoSQP is only a subroutine for solving linear-quadratic programs; a complete method also involves merit function reduction and stepsize selection (here, stepsize refers to that of StoSQP rather than the sketching solver). This paper focuses on uncertainty quantification and statistical inference of online (sketched) Newton methods, which differs significantly from the following literature that focuses on design and convergence of sketched Newton methods. In particular, for many (regression) problems, the objective Hessian can be expressed as $\mathbf{H}=AA^{T}\in\mathbb{R}^{d\times d}$ with a data matrix $A\in\mathbb{R}^{d\times n}$ and $n\geq d$. Then, one can generate a sketch matrix $S\hskip-2.0pt\in\hskip-2.0pt\mathbb{R}^{n\times s}$ and compute the approximate Hessian $\widehat{\mathbf{H}}\hskip-2.0pt=\hskip-2.0ptASS^{T}A^{T}$ilanci2016Iterative developed an iterative Hessian sketch algorithm for solving least-squares problems $\min_{{\bm{x}}}\|A{\bm{x}}-b\|^{2}$ (subject to convex constraints). The authors sketched only the data matrix $A$ rather than both the data matrix $A$ and vector $b$, and established a high-probability convergence result. Lacotte2020Optimal later extended this study by showing the optimal stepsize and convergence rate for Haar sketches. Pilanci2017Newton designed a sketched Newton method that approximates the Hessian using the Johnson–Lindenstrauss transform. Building on this, Agarwal2017Second; Derezinski2019Distributed; Derezinski2020Debiasing; Derezinski2020Precise; Derezinski2021Newton; Lacotte2021Adaptive introduced various sketching methods to explore the trade-off between the computational cost of $\widehat{\mathbf{H}}$ and the convergence rate of the algorithm. In addition to the above series of literature, another type of sketched Newton method is based on Sketch-and-Project framework, where one approximates a generic Hessian inverse $\mathbf{H}^{-1}$ by $S(S^{T}\mathbf{H}S)^{\dagger}S^{T}$ for a sketch matrix $S\in\mathbb{R}^{d\times s}$. See Strohmer2008Randomized; Gower2015Randomized; Luo2016Efficient; Doikov2018Randomized; Gower2019RSN; Derezinski2024Sharp and references therein for the convergence properties of this family of methods.
Compared to deterministic methods for solving Newton systems, such as conjugate gradient and broad preconditioned Krylov (or minimal residual) methods, randomized sketching methods may behave better in terms of improved convergence rates and range of convergence, while requiring less computation and memory (when using suitable sketches) to be scalable and parallelizable (Gower2016Sketch). We refer to Hong2023Constrained for an empirical demonstration of the advantages of sketching solvers over deterministic solvers in the context of SQP methods.
3 Adaptive Inexact StoSQP Method
Let $\mathcal{L}({\bm{x}},{\bm{\lambda}})=f({\bm{x}})+{\bm{\lambda}}^{T}c({\bm{x}})$ be the Lagrangian function of (1.1), where ${\bm{\lambda}}\in\mathbb{R}^{m}$ is the dual vector. Under certain constraint qualifications (introduced later), a necessary condition for $({\bm{x}}^{\star},{\bm{\lambda}}^{\star})$ being a local solution to (1.1) is the KKT conditions: $\nabla\mathcal{L}^{\star}=(\nabla_{{\bm{x}}}\mathcal{L}^{\star},\nabla_{{\bm{\lambda}}}\mathcal{L}^{\star})=({\bm{0}},{\bm{0}})$.
AI-StoSQP applies Newton’s method to the equation $\nabla\mathcal{L}({\bm{x}},{\bm{\lambda}})={\bm{0}}$, involving three steps: estimating the objective gradient and Hessian, (inexactly) solving Newton’s system, and updating the primal-dual iterate. We detail each step as follows. For simplicity, we denote $c_{t}=c({\bm{x}}_{t})$ (similarly, $G_{t}=\nabla c({\bm{x}}_{t})$, $\nabla\mathcal{L}_{t}=\nabla\mathcal{L}({\bm{x}}_{t},{\bm{\lambda}}_{t})$, etc.).
$\bullet$ Step 1: Estimate the gradient and Hessian. We realize a sample $\xi_{t}\sim{\mathcal{P}}$ and estimate the gradient $\nabla f_{t}$ and Hessian $\nabla^{2}f_{t}$ of the objective as
${\bar{g}}_{t}=\nabla F({\bm{x}}_{t};\xi_{t})\quad\quad\text{ and }\quad\quad{\bar{H}}_{t}=\nabla^{2}F({\bm{x}}_{t};\xi_{t}).$
Then, we compute three quantities:
¯${\bar{\nabla}}_{{\bm{x}}}\mathcal{L}_{t}={\bar{g}}_{t}+G_{t}^{T}{\bm{\lambda}}_{t},\quad{\bar{\nabla}}_{{\bm{x}}}^{2}\mathcal{L}_{t}={\bar{H}}_{t}+\sum_{i=1}^{m}({\bm{\lambda}}_{t})_{i}\nabla^{2}c_{i}({\bm{x}}_{t}),\quad B_{t}=\frac{1}{t}\sum_{i=0}^{t-1}{\bar{\nabla}}_{{\bm{x}}}^{2}\mathcal{L}_{i}+\Delta_{t}.$
Here, ¯${\bar{\nabla}}_{{\bm{x}}}\mathcal{L}_{t}$ and ¯${\bar{\nabla}}_{{\bm{x}}}^{2}\mathcal{L}_{t}$ are the estimates of the Lagrangian gradient and Hessian with respect to ${\bm{x}}$, respectively; and $B_{t}$ is a regularized averaged Hessian used in the quadratic program (3.1). We let $\Delta_{t}=\Delta({\bm{x}}_{t},{\bm{\lambda}}_{t})$ be any regularization term ensuring $B_{t}$ to be positive definite in the null space $\{{\bm{x}}\in\mathbb{R}^{d}:G_{t}{\bm{x}}={\bm{0}}\}$. Note that the average $\sum_{i=0}^{t-1}{\bar{\nabla}}_{{\bm{x}}}^{2}\mathcal{L}_{i}/t$ can be updated online. We explain the matrix $B_{t}$ in the following remark.
Remark 3.1.
We note that the Lagrangian Hessian average in $B_{t}$ is over samples $\{\xi_{0},\ldots,\xi_{t-1}\}$, meaning that the Hessian estimate $\bar{H}_{t}$, which depends on the new sample $\xi_{t}$, will only be used in the $(t+1)$-th iteration. Thus, $B_{t}$ and $\Delta_{t}$ are deterministic given $({\bm{x}}_{t},{\bm{\lambda}}_{t})$. In addition, $\Delta_{t}$ can simply be Levenberg-Marquardt type regularization of the form $\delta_{t}I$ with suitably large $\delta_{t}>0$ (Nocedal2006Numerical). The Hessian regularization is standard for nonlinear problems, together with linear independence constraint qualification (LICQ, Assumption 4.1), ensuring that the quadratic program (3.1) is solvable. For convex problems, we just set $\Delta_{t}={\bm{0}}$, $\forall t$. See Bertsekas1982Constrained; Nocedal2006Numerical for various regularization approaches. Moreover, Na2022Hessian showed that Hessian averaging accelerates the local rate of Newton’s method on unconstrained deterministic problems.
$\bullet$ Step 2: Solve the quadratic program. With the above estimates, we solve the quadratic program (QP):
$\min_{{\widetilde{\Delta}}{\bm{x}}_{t}\in\mathbb{R}^{d}}\;\;\frac{1}{2}{\widetilde{\Delta}}{\bm{x}}_{t}^{T}B_{t}\;{\widetilde{\Delta}}{\bm{x}}_{t}+{\bar{g}}_{t}^{T}{\widetilde{\Delta}}{\bm{x}}_{t},\quad\text{s.t.}\;\;c_{t}+G_{t}{\widetilde{\Delta}}{\bm{x}}_{t}={\bm{0}}.$ (3.1)
For the above QP, the objective can be seen as a quadratic approximation of $F({\bm{x}};\xi)$ at $({\bm{x}}_{t};\xi_{t})$, and the constraint can be seen as a linear approximation of $c({\bm{x}})$ at ${\bm{x}}_{t}$. It is easy to observe that solving the above QP is equivalent to solving the following Newton system
$\underbrace{\begin{pmatrix}B_{t}&G_{t}^{T}\\ G_{t}&{\bm{0}}\end{pmatrix}}_{K_{t}}\underbrace{\begin{pmatrix}{\widetilde{\Delta}}{\bm{x}}_{t}\\ {\widetilde{\Delta}}{\bm{\lambda}}_{t}\end{pmatrix}}_{{\widetilde{\bm{z}}}_{t}}=-\underbrace{\begin{pmatrix}{\bar{\nabla}}_{{\bm{x}}}\mathcal{L}_{t}\\ c_{t}\end{pmatrix}}_{{\bar{\nabla}}\mathcal{L}_{t}},$ (3.2)
where $K_{t}$, ¯${\bar{\nabla}}\mathcal{L}_{t}$ are the Lagrangian Hessian and gradient, and ${\widetilde{\bm{z}}}_{t}$ is the exact Newton direction.
Instead of solving the QP (3.1) exactly, we solve it inexactly by an iterative sketching solver. This approach proves more efficient than deterministic solvers, especially when equipped with suitable sketching matrices (Strohmer2008Randomized; Gower2015Randomized; Pilanci2016Iterative; Pilanci2017Newton; Lacotte2020Optimal). In particular, we generate a random sketching matrix $S\in\mathbb{R}^{(d+m)\times s}$, whose column dimension $s\geq 1$ can also be random, and transform the original large-scale linear system to the sketched, small-scale system as
$K_{t}{\bm{z}}_{t}=-{\bar{\nabla}}\mathcal{L}_{t}\hskip 28.45274pt\Longrightarrow\hskip 28.45274ptS^{T}K_{t}{\bm{z}}_{t}=-S^{T}{\bar{\nabla}}\mathcal{L}_{t}.$
Clearly, there are multiple solutions to the sketched system, and ${\bm{z}}_{t}={\widetilde{\bm{z}}}_{t}$ is one of them. We prefer the solution that is closest to the current solution approximation. That is, the $j$-th iteration of the sketching solver has the form (${\bm{z}}_{t,0}={\bm{0}}$)
${\bm{z}}_{t,j+1}=\arg\min_{{\bm{z}}}\|{\bm{z}}-{\bm{z}}_{t,j}\|^{2}\quad\;\;\text{s.t.}\;\;\;\;S_{t,j}^{T}K_{t}{\bm{z}}=-S_{t,j}^{T}{\bar{\nabla}}\mathcal{L}_{t},$ (3.3)
where $S_{t,j}\sim S,\forall j$ are independent and identically distributed and are also independent of $\xi_{t}$. An explicit recursion of (3.3) is given by
${\bm{z}}_{t,j+1}={\bm{z}}_{t,j}-K_{t}S_{t,j}(S_{t,j}^{T}K_{t}^{2}S_{t,j})^{\dagger}S_{t,j}^{T}(K_{t}{\bm{z}}_{t,j}+{\bar{\nabla}}\mathcal{L}_{t}),$ (3.4)
where $(\cdot)^{\dagger}$ denotes the Moore–Penrose pseudoinverse. One can let $s=1$ (i.e., using sketching vectors) so that $S_{t,j}^{T}K_{t}^{2}S_{t,j}$ reduces to a scalar and the pseudoinverse reduces to the reciprocal.
We perform $\tau\geq 1$ iterations of (3.4) and use
$({\bar{\Delta}}{\bm{x}}_{t},{\bar{\Delta}}{\bm{\lambda}}_{t})\coloneqq{\bm{z}}_{t,\tau}$
as the approximate Newton direction. We emphasize that $\tau$ is independent of $t$; thus, we do not require a vanishing approximation error and blow up the computational cost as $t\rightarrow\infty$.
Remark 3.2.
A significant difference between randomized solvers and deterministic solvers is that the approximation error $\|{\bm{z}}_{t,j}-{\widetilde{\bm{z}}}_{t}\|$ of randomized solvers may not be monotonically decreasing as $j$ increases. This subtlety challenges both inference and convergence analysis. In classical optimization world, it is unanimously agreed that if the search direction is asymptotically close to the exact Newton direction (here ${\widetilde{\bm{z}}}_{t}$), then the algorithm will locally behave just like Newton’s method with a similar convergence rate. A precise characterization is called the Dennis-Moré condition (Dennis1974characterization). In our study, although Lemma LABEL:lem:1 shows that the expected error $\mathbb{E}[\|{\bm{z}}_{t,\tau}-{\widetilde{\bm{z}}}_{t}\|\mid{\bm{x}}_{t},\xi_{t}]$ decays exponentially in $\tau$, ${\bm{z}}_{t,\tau}$ can still be far from ${\widetilde{\bm{z}}}_{t}$ for any $t$ and large $\tau$ in the almost sure sense. In particular, (Patel2021Implicit, Theorem 4.2) proved that only a subsequence of $\{\|{\bm{z}}_{t,\tau}-{\widetilde{\bm{z}}}_{t}\|\}_{\tau=0}^{\infty}$ would decrease monotonically with the subsequence indices being also random.
$\bullet$ Step 3: Update the iterate with an adaptive stepsize. With the direction $({\bar{\Delta}}{\bm{x}}_{t},{\bar{\Delta}}{\bm{\lambda}}_{t})={\bm{z}}_{t,\tau}$ from Step 2, we update the iterate $({\bm{x}}_{t},{\bm{\lambda}}_{t})$ by an adaptive stepsize ${\bar{\alpha}}_{t}$:
$({\bm{x}}_{t+1},{\bm{\lambda}}_{t+1})=({\bm{x}}_{t},{\bm{\lambda}}_{t})+{\bar{\alpha}}_{t}\cdot({\bar{\Delta}}{\bm{x}}_{t},{\bar{\Delta}}{\bm{\lambda}}_{t}).$ (3.5)
In principle, the stepsize ${\bar{\alpha}}_{t}$ may rely on the random direction $({\bar{\Delta}}{\bm{x}}_{t},{\bar{\Delta}}{\bm{\lambda}}_{t})$, so it is also random. We allow using any adaptive stepsize selection schemes but require a safeguard condition on ${\bar{\alpha}}_{t}$:
$0<\beta_{t}\leq{\bar{\alpha}}_{t}\leq\eta_{t}\quad\text{ with }\quad\eta_{t}\coloneqq\beta_{t}+\chi_{t},$ (3.6)
where $\{\beta_{t},\eta_{t}\}$ are upper and lower bound sequences and $\chi_{t}$ is the adaptivity gap. We do not require specific stepsize selection schemes beyond the condition (3.6) to achieve our online inference goals. The schemes reported in Berahas2021Sequential; Berahas2023Stochastic; Curtis2021Inexact all adhere to the condition (3.6). See (Berahas2021Sequential, Lemma 3.6) and (Curtis2021Inexact, (25, 28)) for details. Their numerical experiments suggest that adaptive random stepsizes offer promising empirical benefits over non-adaptive deterministic stepsizes (i.e., $\chi_{t}=0$). For sake of completeness, we present a selection scheme from Berahas2021Sequential in Appendix LABEL:app:1.
Algorithm 1 Adaptive Inexact StoSQP Method
1:Input: initial iterate $({\bm{x}}_{0},{\bm{\lambda}}_{0})$, positive sequences $\{\beta_{t},\eta_{t}\}$, an integer $\tau>0$$B_{0}=I$;
2:for $t=0,1,2,\ldots$ do
3:     Realize $\xi_{t}$ and compute ${\bar{g}}_{t}=\nabla f({\bm{x}}_{t};\xi_{t})$, ${\bar{H}}_{t}=\nabla^{2}f({\bm{x}}_{t};\xi_{t})$, and ¯${\bar{\nabla}}_{{\bm{x}}}^{2}\mathcal{L}_{t}$;
4:     Compute the regularized Hessian average $B_{t}=\frac{1}{t}\sum_{i=0}^{t-1}{\bar{\nabla}}_{{\bm{x}}}^{2}\mathcal{L}_{i}+\Delta_{t}$;
5:     Generate sketching matrices $S_{t,j}\sim S$, $\forall j$ iid and iterate (3.4) for $\tau$ times;
6:     Select any adaptive stepsize ${\bar{\alpha}}_{t}$ with $\beta_{t}\leq{\bar{\alpha}}_{t}\leq\eta_{t}$, and update the iterate as (3.5);
7:end for
We combine the above three steps and summarize AI-StoSQP in Algorithm 1. To end this section, we introduce a filtration notation for later use. We let $\mathcal{F}_{t}=\sigma(\{\xi_{i},\{S_{i,j}\}_{j},{\bar{\alpha}}_{i}\}_{i=0}^{t})$, $\forall t\geq 0$ be the $\sigma$-algebra generated by the random variables $\{\xi_{i},\{S_{i,j}\}_{j},{\bar{\alpha}}_{i}\}_{i=0}^{t}$. Moreover, we let $\mathcal{F}_{t-2/3}=\sigma(\{\xi_{i},\{S_{i,j}\}_{j},{\bar{\alpha}}_{i}\}_{i=0}^{t-1}\cup\xi_{t})$, $\mathcal{F}_{t-1/3}=\sigma(\{\xi_{i},\{S_{i,j}\}_{j},{\bar{\alpha}}_{i}\}_{i=0}^{t-1}\cup\xi_{t}\cup\{S_{t,j}\}_{j})$, and have $\mathcal{F}_{t-1}\subseteq\mathcal{F}_{t-2/3}\subseteq\mathcal{F}_{t-1/3}\subseteq\mathcal{F}_{t}$. For consistency, $\mathcal{F}_{-1}$ is the trivial $\sigma$-algebra. With these notation, Algorithm 1 has a generating process as follows: given $({\bm{x}}_{t},{\bm{\lambda}}_{t})$, we first realize $\xi_{t}$ to get the estimates of the gradient ${\bar{g}}_{t}$ and Hessian ${\bar{H}}_{t}$ and derive $\mathcal{F}_{t-2/3}$; then we generate $\{S_{t,j}\}_{j}$ to obtain the inexact Newton direction and derive $\mathcal{F}_{t-1/3}$; then we select the stepsize ${\bar{\alpha}}_{t}$ and derive $\mathcal{F}_{t}$. We also let $(\Delta{\bm{x}}_{t},\Delta{\bm{\lambda}}_{t})$ be the exact Newton direction solved from (3.2) with ¯${\bar{\nabla}}_{{\bm{x}}}\mathcal{L}_{t}$ being replaced by $\nabla_{{\bm{x}}}\mathcal{L}_{t}$.
4 Global Almost Sure Convergence
In this section, we present the global almost sure convergence [^1] Global convergence in nonlinear optimization refers to the convergence to a stationary point from any initialization, in contrast to the convergence to a global solution, which is not achievable without particular problem structures (Nocedal2006Numerical). However, they are equivalent for convex problems as studied for projection-based methods in Duchi2021Asymptotic; Davis2024Asymptotic. for the StoSQP method. We show that the KKT residual $\|\nabla\mathcal{L}_{t}\|$ converges to zero from any initialization. This global convergence serves as a preliminary result of our inference analysis in Section LABEL:sec:4.
We use an adapted augmented Lagrangian function as the Lyapunov function to show the convergence, which has two penalty terms of the form
$\mathcal{L}_{\mu,\nu}({\bm{x}},{\bm{\lambda}})=\mathcal{L}({\bm{x}},{\bm{\lambda}})+\frac{\mu}{2}\|c({\bm{x}})\|^{2}+\frac{\nu}{2}\|\nabla_{{\bm{x}}}\mathcal{L}({\bm{x}},{\bm{\lambda}})\|^{2},\quad\quad\text{with }\mu,\nu>0.$
The first penalty term biases the feasibility error; while, in contrast to standard augmented Lagrangian ($\nu=0$), the extra second penalty term biases the optimality error. We will first show that the inner product between the exact Newton direction $(\Delta{\bm{x}}_{t},\Delta{\bm{\lambda}}_{t})$ and the augmented Lagrangian gradient $\nabla\mathcal{L}_{\mu,\nu}$, with proper parameters $\mu$ and $\nu$, is sufficiently negative (Lemma LABEL:lem:2). Thus, the Newton direction is a descent direction of $\mathcal{L}_{\mu,\nu}$. Then, we will show that the augmented Lagrangian $\mathcal{L}_{\mu,\nu}$ decreases at each step even with an inexact Newton direction (Lemma LABEL:lem:3). This implies that the residual $\|\nabla\mathcal{L}_{t}\|$ finally vanishes to zero (Theorem LABEL:thm:1).
4.1 Assumptions and preliminary results
We state the following assumptions that are standard and proposed in the optimization litera- ture (Kushner1978Stochastic; Bertsekas1982Constrained; Nocedal2006Numerical; Na2022adaptive).
Assumption 4.1.
We assume the existence of a closed, bounded, convex set $\mathcal{X}\times\Lambda$ containing the iterates $\{({\bm{x}}_{t},{\bm{\lambda}}_{t})\}_{t}$, such that $f$ and $c$ are twice continuously differentiable over $\mathcal{X}$. We also assume that the Hessian $\nabla^{2}\mathcal{L}$ is $\Upsilon_{L}$-Lipschitz continuous over $\mathcal{X}\times\Lambda$. In other words,
$\|\nabla^{2}\mathcal{L}({\bm{x}},{\bm{\lambda}})-\nabla^{2}\mathcal{L}({\bm{x}}^{\prime},{\bm{\lambda}}^{\prime})\|\leq\Upsilon_{L}\|({\bm{x}}-{\bm{x}}^{\prime},{\bm{\lambda}}-{\bm{\lambda}}^{\prime})\|,\quad\forall\;({\bm{x}},{\bm{\lambda}}),({\bm{x}}^{\prime},{\bm{\lambda}}^{\prime})\in\mathcal{X}\times\Lambda.$ (4.1)
Furthermore, we assume that the constraints Jacobian $G_{t}$ has full row rank with $G_{t}G_{t}^{T}\succeq\gamma_{G}I$ for a constant $\gamma_{G}>0$. Additionally, the regularization $\Delta_{t}$ ensures that $B_{t}$ satisfies $\|B_{t}\|\leq\Upsilon_{B}$ and ${\bm{x}}^{T}B_{t}{\bm{x}}\geq\gamma_{RH}\|{\bm{x}}\|^{2}$ for any ${\bm{x}}\in\{{\bm{x}}\in\mathbb{R}^{d}:G_{t}{\bm{x}}={\bm{0}}\}$, for some constants $\gamma_{RH},\Upsilon_{B}>0$.
Assumption 4.1 assumes $G_{t}$ has full row rank, which is referred to as the linear independence constraint qualification (LICQ). LICQ is a common constraint qualification ensuring the uniqueness of the dual solution, and is also necessary for the inference analysis (see (Duchi2021Asymptotic, Assumption B) and (Davis2024Asymptotic, Example 2.1)). LICQ and conditions on $B_{t}$ are critical for SQP methods; they imply QP (3.1) has a unique solution (Nocedal2006Numerical, Lemma 16.1). The Lipschitz continuity of the Hessian matrix is also standard for analyzing Newton’s method.
The bounded iterates condition is commonly assumed in the literature on (stochastic) nonlinear nonconvex optimization, both for first-order gradient-based methods (Bolte2013Proximal; Song2014Weak; Davis2016sound; Davis2019Stochastic; Atchade2017perturbed; Asi2019Stochastic; Liu2023Statistical) and for second-order SQP methods ((Bertsekas1982Constrained, Proposition 4.15), (Nocedal2006Numerical, Theorem 18.3)). This assumption ensures that all functions with their gradients and Hessians are bounded over $\mathcal{X}\times\Lambda$ as long as they are smooth. Some literature replaces the bounded iterates condition by directly imposing boundedness on the gradients and Hessians of the objective and constraints, although the main use of the condition in the proof is rather similar (Berahas2021Sequential; Berahas2023Stochastic; Curtis2021Inexact; Ramprasad2022Online; Liu2023Online).
We provide two justifications for the boundedness condition. First, in our study, the StoSQP iterates presumably track a deterministic feasible set $\{{\bm{x}}\in\mathbb{R}^{d}:c({\bm{x}})={\bm{0}}\}$, so we believe that an unbounded iteration sequence is generally rare especially when the feasible set is bounded. Second, a practical way to enforce the boundedness condition may be through adaptive truncation (Andrieu2005Stability; Liang2010Trajectory). Under some conditions on the Markov transition kernel of the iteration sequence, one can show that the truncation occurs only finitely many times, ensuring that the convergence and asymptotic behavior are finally not affected by the truncation.
We also impose bounded moment conditions on the stochastic estimates ${\bar{g}}_{t}$ and ${\bar{H}}_{t}$.
Assumption 4.2.
We assume $\mathbb{E}[{\bar{g}}_{t}\mid{\bm{x}}_{t}]=\nabla f_{t}$, $\mathbb{E}[{\bar{H}}_{t}\mid{\bm{x}}_{t}]=\nabla^{2}f_{t}$, and assume the following moment conditions when needed: for a constant $\Upsilon_{m}>0$,
[…]